# DiffGAN: A Test Generation Approach for Differential Testing of Deep Neural Networks

### v2

@@ 4.2 Baseline Approach
Based on the first criterion we identified DiffChaser [8], BET [11], Diverget [6], DRfuzz [10], and DFlare [9], as candidates. However, BET and Diverget were excluded due to their lack of open-source implementations. We also decided not to include DiffChaser in our experiments, as DRfuzz has demonstrated superior performance in generating a greater number of trigger inputs [10]. Consequently, we selected DRfuzz as a baseline to evaluate the effectiveness of our proposed approach. We should note that DFlare was excluded because its test generation process is inherently constrained by the number of initial test inputs. The method applies multiple mutation rules to each seed input but halts once a triggering input is found, preventing further mutation. As a result, DFlare can generate at most one triggering input per initial seed, significantly limiting the overall number of triggering inputs. To mitigate this limitation, we attempted to repeatedly feed the same seed inputs into DFlare to increase the number of generated inputs. However, due to DFlare’s low randomness in input generation, repeated executions produced highly redundant outputs. This redundancy is further amplified by the limited number of mutation operations and the exploitation of the same mutation operators that successfully increase behavioral disagreements between the models under test over the generations. As a result, when repeatedly fed with seed inputs, DFlare’s mutation process introduces limited variations, failing to produce a diverse set of triggering inputs. To quantify this limitation, we compared the diversity of the inputs generated by DFlare, DRFuzz, and our approach. Our results showed that DFlare’s generated inputs had significantly lower diversity scores than those produced by both DRFuzz and DiffGAN. Therefore, we excluded this baseline in our study.
@@ 5 Empirical Results
Table IV: RQ1 results: Number of generated triggering images for models trained on MNIST dataset using DiffGAN and DRfuzz
@@
Table V: RQ1 results: Number of generated triggering images for models trained on Cifar-10 dataset using DiffGAN and DRfuzz
@@
To systematically compare DiffGAN and DRfuzz in terms of generating valid triggering inputs within a given time budget, we considered the inputs generated in RQ1 by both methods. More specifically, we considered all four model pairs using the MNIST dataset and ran both approaches for two hours on these models. We chose to focus on MNIST in this RQ due to how resource-intensive is the evaluation process. Indeed, validity assessments require significant manual effort, particularly for labeling generated inputs to ensure they are contextually valid. The lower resolution and greater complexity of CIFAR-10 images further complicate this process, as accurately labeling and validating inputs in CIFAR-10 is considerably more challenging compared to MNIST. Given these challenges, we opted to limit our validity study to the MNIST dataset to ensure its feasibility.
@@
Selecting representative samples. Due to the large number of triggering inputs generated by DiffGAN and DRfuzz, we selected a sample of triggering inputs from each aggregated dataset. For each setting, we determined the sample size that can confidently represent the whole corresponding dataset with a 95% confidence level and a margin of error of 5% [^5] As computed from https://www.calculator.net/sample-size-calculator.html. As a result, in our experiments, the sample size varied from 361 to 381 across the four settings that we considered, as shown in Table VIII. This approach helped choose a minimal subset that was large enough to provide results that accurately reflected the overall distribution of inputs generated by DiffGAN and DRfuzz [62].
Manual validity checking. From the non-redundant set, we randomly sampled images for validation, according to the sample size determined the previous step. To ensure unbiased evaluation, we shuffled images generated by both DiffGAN and DRfuzz and presented them anonymously to three independent expert evaluators. These experts were tasked with assessing the validity of each image based on predefined criteria. Specifically, they labeled images as either valid or invalid. Invalid images were defined as those that displayed two numbers, no number, or other issues such as (1) Images that were a 50/50 blend of two digits and could not be clearly identified as a specific digit, (2) Incomplete or blended digits that could not be classified into one category by consensus, (3) Images containing random shapes or non-digit drawings, or (4) Images that were completely white or dark, with no discernible content. Examples of valid and invalid inputs generated by both approaches are depicted in Figure 4. To determine the validity of each image, we considered the agreement among the evaluators. If at least two out of three evaluators labeled an image as valid, we considered it valid. To quantify the level of agreement between the evaluators, we employ Cohen’s kappa [63]—a statistical measure that accounts for agreement occurring by chance. In our study, the value of Cohen’s kappa is 70%, indicating a high level of agreement among the evaluators and underscoring the reliability of the manual labeling process.
@@
Finally, we compared the expected number of valid inputs generated by DiffGAN and DRfuzz to assess which method was more effective at producing valid triggering inputs within the same time budget of two hours. The results presented in Table VIII, clearly suggest that DiffGAN significantly outperforms DRfuzz in generating more valid triggering inputs. For instance, in the first setting, it is estimated that DiffGAN generated 8,292 valid triggering inputs, whereas DRfuzz produced only 1,999 such inputs. This corresponds to a validity ratio of 91.25% for DiffGAN compared to 82.46% for DRfuzz. Importantly, those results were consistent across all the settings we considered. To further analyze the statistical significance of our results we performed a statistical analysis using the Wilcoxon signed-rank test [65], with a significance level of $\alpha=0.05$, to assess whether DiffGAN significantly outperforms DRfuzz in generating valid triggering inputs. The Wilcoxon signed-rank test, a non-parametric method, is used to compare the medians of continuous variables for paired samples and is ideal for our case since it does not require any assumptions about the data distribution. Instead of considering the number of valid inputs, we relied on the validity ratio due to differences in sample sizes. We evaluated the validity ratios from 4000 samples for DiffGAN and 4000 samples for DRfuzz, selected via bootstrapping across four different scenarios (1000 samples per scenario). The Wilcoxon signed-rank test, with w-statistic = 343465.5 and a p-value of 0.000001, confirmed that DiffGAN significantly outperformed DRfuzz at generating valid triggering inputs across all scenarios.
Table VIII: Validity Comparison
@@
The results, as shown in Tables IX and X, consistently demonstrate that DiffGAN outperforms DRfuzz in generating more diverse triggering inputs, when considering both metrics, highlighting its effectiveness in exploring a broader range of input variations.
@@
Table IX: Diversity scores of MNIST
@@
Table X: Diversity scores of Cifar-10
@@
This improved accuracy can be attributed to several factors. DiffGAN leverages the combination of GAN and NSGA-II to generate highly diverse triggering inputs. Unlike DRfuzz, which applies mutation rules to existing seed inputs, DiffGAN explores a significantly broader latent input space, enabling the generation of more diverse inputs that expose more pronounced behavioral differences between the models. These behavioral differences are crucial for the RF model to learn the conditions when each classifier performs better. Additionally, DiffGAN optimizes its inputs to both maximize diversity and reveal discrepancies, which also enables the RF model to develop a more accurate understanding of the conditions under which one classifier outperforms the other.
Though the accuracy of RF models can still be improved, our main goal here was to show the usefulness of DiffGAN for one important application, not to optimize model selection. The latter would require the labeling of a much larger dataset and possibly the fine-tuning of the RF models or experimenting with other machine learning techniques, which is out of the scope of this paper.
@@ 6 Discussions
Practical Applicability: The black-box nature of DiffGAN makes it agnostic to the models under test, as it does not require any assumptions on their architectural similarity. This flexibility allows DiffGAN to be applied across diverse practical scenarios where models with similar accuracy levels are developed independently or trained using different data and optimization techniques. The ability of DiffGAN to generate diverse and valid triggering inputs makes it a practical approach for evaluating and comparing DNN models with similar accuracy levels. DiffGAN is therefore particularly valuable for model selection, regression testing, and model compression, that are common application scenarios in ML development pipelines. In model selection, the triggering inputs generated by DiffGAN can help identify behavioral discrepancies enabling fine-grained analysis beyond accuracy metrics, ensuring the selection of the most adequate model. In regression testing, DiffGAN can be used to generate cases where an updated model underperforms compared to its predecessor. In model compression validation, DiffGAN can be used to understand when compressed models fare less well than their original versions and whether this is acceptable. Another important application of DiffGAN is ML-based model selection, a scenario informed by our industrial partner. In some autonomous vehicles, classifiers with similar accuracy levels are dynamically selected on the fly to classify sensor-captured images. Since different classifiers specialize in distinct regions of the feature space, selecting the most suitable one for each input improves reliability. Additionally, certain classifiers require more computational resources than others, making dynamic selection crucial for optimizing inference time and energy consumption, particularly in embedded systems with limited resources. By generating diverse inputs that expose model discrepancies, DiffGAN enhances the accuracy of model selection mechanisms since their training dataset is enriched by such inputs and more informative. The relevance of this application is further supported by existing literature on ML-based model selection [51, 52].
@@
Generalisability: Although DiffGAN is primarily designed for image classification tasks, our approach can be extended to other domains such as image regression and segmentation. This extension can be achieved by adapting the divergence fitness function. For instance, in regression tasks, triggering inputs can be defined as those where the difference between the models’ outputs exceeds a predefined threshold. The divergence fitness function can then be tailored to capture a meaningful divergence metric, such as absolute or relative differences between regression model outputs, depending on the specific context. The primary objective would be to maximize the discrepancy between numerical outputs. For segmentation tasks, triggering inputs can be determined based on differences in predicted labels, variations in segmented shapes, or discrepancies in segmented regions. Once a triggering input is defined, the divergence fitness function can be customized to highlight these differences in the models’ outputs effectively. We should note that, for both regression and segmentation, the diversity-based fitness function will remain unchanged as long as the inputs are images. However, if the input type varies, the fitness function should be adapted accordingly to ensure an appropriate measure of diversity. We should also note that DiffGAN is not applicable to tasks involving multi-modal inputs or time-series data in its current form. This arises from the fact that our framework relies on GANs, which are specifically designed for image generation and manipulation. Supporting multi-modal or time-series tasks would require several modifications, such as integrating alternative generative architectures (e.g., Variational Autoencoders or Recurrent Neural Networks) and designing new strategies for defining and optimizing triggering inputs. While DiffGAN is designed for image classification, its black-box differential testing approach can be extended to LLMs by adapting input generation using trained text-based models, employing embedding similarity or log-likelihood divergence as a fitness function, and integrating language model discriminators for validity filtering. As future work, we aim to extend DiffGAN to support a broader range of ML tasks beyond image classification. This includes adapting the fitness function to better capture model discrepancies and the diversity of the generated inputs. We also aim to explore how DiffGAN can be adapted for other types of inputs, such as textual and audio inputs.
GAN Mode Collapse Mitigation: One of the key challenges of GANs is their propensity for mode collapse, which limits the diversity of the generated inputs. To mitigate this issue, our approach integrates multiple strategies during both the training phase and test input generation to improve the generation of diverse inputs. Specifically, during training, we ensured that GANs were properly optimized, achieving low FID and KID scores (Section 3.2). These metrics, widely recognized for evaluating the quality and diversity of generated samples, were used to confirm that the trained GANs effectively captured the diversity of the underlying training data distribution [33]. Low FID and KID scores indicate that the GAN not only avoids mode collapse but also generates realistic and diverse samples. Another factor we used to improve the diversity of the data was generating an augmented training dataset to train the GAN by applying different image transformations (Section 3.1). This data augmentation exposed the GAN to a more diversified training dataset, mitigating the risk of mode collapse. Furthermore, during the test case generation process, we explicitly incorporated diversity as a key fitness function. As detailed in Section 3.3.2, we clustered generated images based on their features and ensured that new images were distinct from previously generated ones. This approach guided the GAN to explore a broader range of outputs, reducing the likelihood of mode collapse and maintaining high diversity in the generated triggering inputs. Finally, our evaluation results demonstrated that our approach consistently outperformed the baseline in generating more diverse triggering inputs, thus validating the effectiveness of our diversity-enhancing strategies.
@@
As shown in Figure 6, we found that DiffGAN scales linearly with increasing dataset size, with a training time ranging from 371 to 607 seconds. We note that this increase remains practically acceptable, as GAN training is a one-time process per dataset, and differential testing is neither frequent nor a real-time task. Thus, training costs do not significantly impact the feasibility of using DiffGAN for large-scale differential testing. Additionally, it is important to clarify that DiffGAN does not strictly require training a GAN from scratch, as it treats GANs (or any other generative model) as black-box components. This means that, for large-scale datasets such as ImageNet, DiffGAN can leverage pretrained GANs or generators that have already been trained on such datasets [74, 75], significantly reducing the computational burden while maintaining its ability to generate diverse and valid test cases. Existing GAN models trained on large datasets can be seamlessly integrated into DiffGAN without requiring additional training, making our approach even more scalable and applicable beyond the datasets tested in our current study.

### v3

@@ 4.2 Baseline Approach
Based on the first criterion we identified DiffChaser [8], BET [11], Diverget [6], DRfuzz [10], and DFlare [9], as candidates. However, BET and Diverget were excluded due to their lack of open-source implementations. We also decided not to include DiffChaser in our experiments, as DRfuzz has demonstrated superior performance in generating a greater number of trigger inputs [10]. Consequently, we selected DRfuzz as a baseline to evaluate the effectiveness of our proposed approach. We should note that DFlare was excluded because its test generation process is inherently constrained by the number of initial test inputs. The method applies multiple mutation rules to each seed input but halts once a triggering input is found, preventing further mutation. As a result, DFlare can generate at most one triggering input per initial seed, significantly limiting the overall number of triggering inputs. To mitigate this limitation, we attempted to repeatedly feed the same seed inputs into DFlare to increase the number of generated inputs. However, due to DFlare’s low randomness in input generation, repeated executions produced highly redundant outputs. This redundancy is further amplified by the limited number of mutation operations and the exploitation of the same mutation operators that successfully increase behavioral disagreements between the models under test over the generations. As a result, when repeatedly fed with seed inputs, DFlare’s mutation process introduces limited variations, failing to produce a diverse set of triggering inputs. To quantify this limitation, we compared the diversity of the inputs generated by DFlare, DRFuzz, and our approach. Our results showed that DFlare’s generated inputs had significantly lower diversity scores than those produced by both DRFuzz and DiffGAN. Therefore, we excluded this baseline in our study. Detailed results of this comparison are available in Appendix 1.
@@ 5 Empirical Results
Table IV:  RQ1 results: Number of generated triggering images for models trained on MNIST dataset using DiffGAN and DRfuzz.The #TI (Initial) column indicates the number of triggering inputs identified in the original test dataset for each set of models.
@@
Table V: RQ1 results: Number of generated triggering images for models trained on Cifar-10 dataset using DiffGAN and DRfuzz. The #TI (Initial) column indicates the number of triggering inputs identified in the original test dataset for each set of models.
@@
proach):
To systematically compare DiffGAN and DRfuzz in terms of generating valid triggering inputs within a given time budget, we considered the inputs generated in RQ1 by both methods. More specifically, we considered all four model pairs using the MNIST dataset and two model pairs for the CIFAR-10 dataset. We ran both approaches for two hours on these models. We did not consider all model pairs in this RQ as the validity assessments require significant manual effort, particularly for labeling generated inputs to ensure they are contextually valid. The lower resolution and greater complexity of CIFAR-10 images further complicate this process, as accurately labeling and validating inputs in CIFAR-10 is considerably more challenging compared to MNIST. Given these challenges, we opted to limit our validity study to the MNIST dataset to ensure its feasibility.
@@
Selecting representative samples. Due to the large number of triggering inputs generated by DiffGAN and DRfuzz, we selected a sample of triggering inputs from each aggregated dataset. For each setting, we determined the sample size that can confidently represent the whole corresponding dataset with a 95% confidence level and a margin of error of 5% [^5] As computed from https://www.calculator.net/sample-size-calculator.html. As a result, in our experiments, the sample size varied from 364 to 384 across the six settings that we considered, as shown in Tables VIII and IX. This approach helped choose a minimal subset that was large enough to provide results that accurately reflected the overall distribution of inputs generated by DiffGAN and DRfuzz [62].
Manual validity checking. From the non-redundant set, we randomly sampled images for validation, according to the sample size determined the previous step. To ensure unbiased evaluation, we shuffled images generated by both DiffGAN and DRfuzz and presented them anonymously to three independent expert evaluators. These experts were tasked with assessing the validity of each image based on predefined criteria. Specifically, they labeled images as either valid or invalid. For example for MNIST dataset, invalid images were defined as those that displayed two numbers, no number, or other issues such as (1) Images that were a 50/50 blend of two digits and could not be clearly identified as a specific digit, (2) Incomplete or blended digits that could not be classified into one category by consensus, (3) Images containing random shapes or non-digit drawings, or (4) Images that were completely white or dark, with no discernible content. Examples of valid and invalid inputs generated by both approaches are depicted in Figure 4. To determine the validity of each image, we considered the agreement among the evaluators. If at least two out of three evaluators labeled an image as valid, we considered it valid. To quantify the level of agreement between the evaluators, we employ Cohen’s kappa [63]—a statistical measure that accounts for agreement occurring by chance. In our study, the value of Cohen’s kappa is 70%, indicating a high level of agreement among the evaluators and underscoring the reliability of the manual labeling process.
@@
Finally, we compared the expected number of valid inputs generated by DiffGAN and DRfuzz to assess which method was more effective at producing valid triggering inputs within the same time budget of two hours. The results, presented in Tables VIII and IX , clearly suggest that DiffGAN significantly and consistently outperforms DRfuzz in generating more valid triggering inputs for all the model pairs and datasets we considered in our experiments. For instance, in the first setting, it is estimated that DiffGAN generated 8,292 valid triggering inputs, whereas DRfuzz produced only 1,999 such inputs. This corresponds to a validity ratio of 91.25% for DiffGAN compared to 82.46% for DRfuzz. To further analyze the statistical significance of our results we performed a statistical analysis using the Wilcoxon signed-rank test [65], with a significance level of $\alpha=0.05$, to assess whether DiffGAN significantly outperforms DRfuzz in generating valid triggering inputs. The Wilcoxon signed-rank test, a non-parametric method, is used to compare the medians of continuous variables for paired samples and is ideal for our case since it does not require any assumptions about the data distribution. Instead of considering the number of valid inputs, we relied on the validity ratio due to differences in sample sizes. We evaluated the validity ratios from 4000 samples for DiffGAN and 4000 samples for DRfuzz, selected via bootstrapping across four different scenarios (1000 samples per scenario). The Wilcoxon signed-rank test, with w-statistic = 343465.5 and a p-value of 0.000001, confirmed that DiffGAN significantly outperformed DRfuzz at generating valid triggering inputs across all scenarios.
Table VIII: Validity Comparison MNIST
@@
Table IX: Validity Comparison Cifar-10
M1, M2 DiffGAN 384 0.78 [0.7353, 0.8196] 113,393
DRFuzz 382 0.67 [0.6200, 0.7111] 31,422
M3, M4 DiffGAN 384 0.71 [0.6619, 0.7531] 79,613
DRFuzz 378 0.58 [0.5293, 0.6286] 7,782
@@
The results, as shown in Tables X and XI, consistently demonstrate that DiffGAN outperforms DRfuzz in generating more diverse triggering inputs, when considering both metrics, highlighting its effectiveness in exploring a broader range of input variations.
@@
Table X: Diversity scores of MNIST
@@
Table XI: Diversity scores of Cifar-10
@@
Table XII: Categorization of Model Behavior on 1,000 Triggering Inputs per Model Pair (Counts with Percentages)
Models Method Only Model A is correct Only Model B is correct Both models are incorrect
M1, M2 DRfuzz 282 (28.2%) 480 (48.0%) 238 (23.8%)
DiffGAN 121 (12.1%) 742 (74.2%) 137 (13.7%)
M3, M4 DRfuzz 294 (29.4%) 586 (58.6%) 120 (12.0%)
DiffGAN 285 (28.5%) 620 (62.0%) 95 (9.5%)
M5, M6 DRfuzz 267 (26.7%) 560 (56.0%) 173 (17.3%)
DiffGAN 198 (19.8%) 687 (68.7%) 115 (11.5%)
M7, M8 DRfuzz 303 (30.3%) 569 (56.9%) 128 (12.8%)
DiffGAN 298 (29.8%) 592 (59.2%) 110 (11.0%)
Model A refers to the first model in each pair (e.g., M1, M3, etc.).
@@
We also manually analyzed the types of triggering inputs generated by DiffGAN and DRFuzz by randomly sampling 1,000 labeled valid triggering images for each method and model pair. Each input was categorized into one of the following types: (1) only Model A produces the correct label, (2) only Model B produces the correct label, and (3) both models produce incorrect but different labels. Here, Model A refers to the first model in each pair, and Model B to the second. This categorization enabled us to quantify the nature of disagreements revealed by the generated inputs. As shown in Table XII, both DiffGAN and DRFuzz exhibit all disagreement types, however, in different proportions. Indeed, DiffGAN generates significantly fewer inputs where both models are incorrect. This is practically important as we want to generate as many inputs that distinguish the two models’ behavior as possible, regarding conditions when they can be trusted, in order to learn which model to trust for a given input. In turn, this is useful to devise accurate model selectors, for example, at run-time. In conclusion, these results further explain why DiffGAN yields better model selectors than DRFuzz.
@@
This improved accuracy can be attributed to several factors. DiffGAN leverages the combination of GAN and NSGA-II to generate highly diverse triggering inputs. Unlike DRfuzz, which applies mutation rules to existing seed inputs, DiffGAN explores a significantly broader latent input space, enabling the generation of more diverse inputs that expose more pronounced behavioral differences between the models. These behavioral differences are crucial for the RF model to learn the conditions when each classifier performs better. Additionally, DiffGAN optimizes its inputs to both maximize diversity and reveal discrepancies, which also enables the RF model to develop a more accurate understanding of the conditions under which one classifier outperforms the other. Though the accuracy of RF models can still be improved, our main goal here was to show the usefulness of DiffGAN for one important application, not to optimize model selection. The latter would require the labeling of a much larger dataset and possibly the fine-tuning of the RF models or experimenting with other machine learning techniques, which is out of the scope of this paper.
@@ 6 Discussions
Practical Applicability: The black-box nature of DiffGAN makes it agnostic to the models under test, as it does not require any assumptions on their architectural similarity. This flexibility allows DiffGAN to be applied across diverse practical scenarios where models with similar accuracy levels are developed independently or trained using different data and optimization techniques. The ability of DiffGAN to generate diverse and valid triggering inputs makes it a practical approach for evaluating and comparing DNN models with similar accuracy levels. DiffGAN is therefore particularly valuable for model selection, regression testing, and model compression, which are common application scenarios in ML development pipelines. In model selection, the triggering inputs generated by DiffGAN can help identify behavioral discrepancies enabling fine-grained analysis beyond accuracy metrics, ensuring the selection of the most adequate model. In regression testing, DiffGAN can be used to generate cases where an updated model underperforms compared to its predecessor. In model compression validation, DiffGAN can be used to understand when compressed models fare less well than their original versions and whether this is acceptable. Another important application of DiffGAN is ML-based model selection, a scenario informed by our industrial partner. In some autonomous vehicles, classifiers with similar accuracy levels are dynamically selected on the fly to classify sensor-captured images. Since different classifiers specialize in distinct regions of the feature space, selecting the most suitable one for each input improves reliability. Additionally, certain classifiers require more computational resources than others, making dynamic selection crucial for optimizing inference time and energy consumption, particularly in embedded systems with limited resources. By generating diverse inputs that expose model discrepancies, DiffGAN enhances the accuracy of model selection mechanisms since their training dataset is enriched by such inputs and is more informative. The relevance of this application is further supported by existing literature on ML-based model selection [51, 52].
@@
Generalisability: Although DiffGAN is primarily designed for image classification tasks, our approach can be extended to other domains such as image regression and segmentation. This extension can be achieved by adapting the divergence fitness function. For instance, in regression tasks, triggering inputs can be defined as those where the difference between the models’ outputs exceeds a predefined threshold. The divergence fitness function can then be tailored to capture a meaningful divergence metric, such as absolute or relative differences between regression model outputs, depending on the specific context. The primary objective would be to maximize the discrepancy between numerical outputs. For segmentation tasks, triggering inputs can be determined based on differences in predicted labels, variations in segmented shapes, or discrepancies in segmented regions. Once a triggering input is defined, the divergence fitness function can be customized to highlight these differences in the models’ outputs effectively. We should note that, for both regression and segmentation, the diversity-based fitness function will remain unchanged as long as the inputs are images. However, if the input type varies, the fitness function should be adapted accordingly to ensure an appropriate measure of diversity. We should also note that DiffGAN is not applicable to tasks involving multi-modal inputs or time-series data in its current form. This arises from the fact that our framework relies on GANs, which are specifically designed for image generation and manipulation. Supporting multi-modal or time-series tasks would require several modifications, such as integrating alternative generative architectures (e.g., Variational Autoencoders or Recurrent Neural Networks) and designing new strategies for defining and optimizing triggering inputs. While DiffGAN is designed for image classification, its black-box differential testing approach can be extended to LLMs by adapting input generation using trained text-based models, employing embedding similarity or log-likelihood divergence as a fitness function, and integrating language model discriminators for validity filtering. As future work, we aim to extend DiffGAN to support a broader range of ML tasks beyond image classification. This includes adapting the fitness function to capture better model discrepancies and the diversity of the generated inputs. We also aim to explore how DiffGAN can be adapted for other types of inputs, such as textual and audio inputs.
Furthermore, our evaluation includes both highly accurate model pairs (with accuracy above 98%) and relatively less accurate ones (ranging from 80% to 97%) to enhance the generalizability of our results. For highly accurate models, the generation of additional triggering inputs is considerably more challenging due to the scarcity of disagreement regions in the input space. For less accurate models, while behavioral disagreements may be more frequent, the key challenge remains in generating inputs that are not only valid but also diverse, so as to avoid trivial or uninformative disagreements. As shown in our ablation study (Table III), maximizing only divergence without accounting for diversity leads to large numbers of triggering inputs, but many are similar, offering limited insight into the boundary behaviors of the models. While the original test dataset contains triggering inputs for each model pair, these inputs could be redundant and typically do not sufficiently capture the diverse conditions under which behavioral differences between models occur. Relying solely on these inputs would not provide a comprehensive view of how and when models diverge in their predictions. DiffGAN addresses these challenges by generating, on average, four times more triggering inputs than the SOTA baseline, while also significantly improving both the validity and the diversity of these inputs. This richer and more diverse set of valid triggering inputs enables a more comprehensive behavioral analysis of model pairs.
GAN Mode Collapse Mitigation: One of the key challenges of GANs is their propensity for mode collapse, which limits the diversity of the generated inputs. To mitigate this issue, our approach integrates multiple strategies during both the training phase and test input generation to promote diverse inputs. Specifically, to detect and monitor instability during training, we systematically tracked both the generator and discriminator loss curves throughout the training process. This allowed us to observe convergence behavior and identify any signs of instability, such as divergence or mode collapse. In addition, as detailed in Section 3.2, we employed the Optuna framework to automate hyperparameter optimization. Optuna was used to tune learning rates, batch sizes, number of epochs, and other key model settings, to minimize FID and KID scores, metrics widely accepted for evaluating both the quality and diversity of generated outputs. We further enhanced the robustness of this process by averaging FID and KID scores across ten independent runs per configuration. Optuna’s early-stopping (pruner) mechanism helped discard unstable or underperforming configurations, contributing to a more stable and efficient optimization process. These measures allowed us to identify configurations that led to more stable and better-quality training outcomes. The final trained models consistently achieved low FID and KID scores, which indicate strong convergence and the ability to capture the underlying data distribution without collapsing to limited modes (Section 3.2).
We incorporated further strategies to explicitly enhance output diversity and reduce the likelihood of mode collapse. These included both modifications to the GAN training process and techniques applied during test input generation and evaluation:
(1) Data augmentation during GAN training: We applied a variety of image transformations to augment the original training dataset. This strategy exposed the GAN to a more heterogeneous input distribution, which is essential for preventing overfitting to narrow patterns and mitigating the risk of mode collapse. By training on this enriched dataset, the GAN learned to generalize better and produce outputs that reflect a wider range of valid data characteristics.
(2) Diversity-guided test input generation using feature-based clustering: During the generation of test inputs, we explicitly incorporated diversity as a key part of the search objective. By clustering generated images based on their feature representations, we ensured that newly generated samples were not only valid but also meaningfully different from previously generated ones. This clustering-based strategy prevented the GAN from generating repetitive or redundant outputs. It encouraged broader exploration of the latent space, thereby minimizing the risk of mode collapse and mitigating its potential consequences during input synthesis.
(3) Quantitative diversity evaluation using two distinct metrics: To assess the diversity of the outputs objectively, we used two separate diversity metrics and compared our method against multiple baselines. These metrics evaluated both the distributional spread and uniqueness of the generated samples. Our method consistently outperformed the baselines, confirming that our strategies led to more diverse and comprehensive coverage of the input space. This empirical evidence further supports the claim that our model avoided collapse to limited modes.
In combination, these strategies formed a multi-stage framework to avoid mode collapse and ensure high-quality, diverse outputs. Each component reinforced the others: enriched data improved variation, clustering promoted diverse generation, Optuna-based tuning with loss monitoring and early stopping ensured stability, and diversity metrics validated effectiveness. This integrated process not only reduced the likelihood of mode collapse but also prevented its potential consequences by identifying and correcting instability early and ensuring that the diversity and realism of the generated outputs were preserved.
@@
As shown in Figure 6, we found that DiffGAN scales linearly with increasing dataset size, with a training time ranging from 371 to 607 seconds. We note that this increase remains practically acceptable, as GAN training is a one-time process per dataset, and differential testing is neither frequent nor a real-time task. Thus, training costs do not significantly impact the feasibility of using DiffGAN for large-scale differential testing. Additionally, it is important to clarify that DiffGAN does not strictly require training a GAN from scratch, as it treats GANs (or any other generative model) as black-box components. This means that, for large-scale datasets such as ImageNet, DiffGAN can leverage pretrained GANs or generators that have already been trained on such datasets [74, 75]. This would reduce the computational burden of training new GANs while maintaining the ability of DiffGAN to generate diverse and valid test cases. We consider incorporating pretrained GANs and evaluating their effect on triggering input quality and computational savings as future work.